OpenStudy (anonymous):

Solve, please explain. h - 8/3 = 1/8

OpenStudy (mertsj):

Add 8/3 to both sides.

OpenStudy (hoblos):

h= 1/8 +8/3 h= 3/24 + 64/24 = 67/24

OpenStudy (anonymous):

It's : H - 8 over 3 = 1 over 8. So Hoblos, is your answer still correct?

OpenStudy (mertsj):

\[\frac{h-8}{3}=\frac{1}{8}\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

Yup/

OpenStudy (mertsj):

Then cross multiply

OpenStudy (mertsj):

8h-64=3

OpenStudy (mertsj):

Add 64 to both sides: 8h = 67

OpenStudy (mertsj):

Divide by 8: h = 67/8

OpenStudy (hoblos):

nope it will change (h-8)/3 = 1/8 h-8= 3/8 h= 3/8 + 64/8 h= 67/8

OpenStudy (anonymous):

Thank you!